# Automated Label Unification for Multi-Dataset Semantic Segmentation with GNNs

### v1

@@ 4.3 Ablation Study
To further explore the ability of our GNNs to construct a unified label space, we compared our method with three different approaches for constructing a unified label space. Variant 1 concatenates all dataset label spaces as the unified space $\bigcup_{i=1}^{D}L_{i}$. Variant 2 constructs a unified label space by clustering text features based on cosine similarity using the DBSCAN [7] algorithm. Variant 3 conducts an ablation experiment by removing label description module. The experimental results are shown in the Table 5. The results indicate that the label space constructed based on GNNs can better assist in the learning of segmentation models. Our approach does not waste model capacity on dataset recognition compared to variant 1. Moreover, compared to variant 2, our method can handle classes with the same name but different granularity. By leveraging label descriptions to provide rich semantic context information, our method could construct a better label space.
@@ 5 Conclusion
Limitations. While our method offers advantages in multi-dataset training, there are limitations to consider. For instance, the model’s performance may decrease when a single batch cannot contain samples from all datasets, imposing significant GPU memory requirements during training. Additionally, challenges related to domain shift, which were not discussed in this paper, may affect the model’s performance on different datasets.
@@ Appendix A Training Strategy
The training process is divided into three distinct stages: Multi-SegHead Training Stage, Alternating GNNs Training Stage and Alternating SegNet Training Stage, as shown in Algorithm 1. Initially, we commence with the Multi-SegHead Training Stage for 100k iterations. Subsequently, we alternate between the Alternating GNNs Training Stage and Alternating SegNet Training Stage for a total of three cycles, each lasting 20k iterations. Notably, we prioritize the Alternating GNNs Training stage at the beginning of each cycle.
@@
$\mathcal{L}_{m}(\mathbf{y},\mathbf{s})=\sum^{L}_{k=1}y_{k}\log(\operatorname{softmax}(\mathbf{W}_{i}\textbf{p})_{k}).$ (9)
@@ Appendix B Selection of Unified Label Node Quantity and Initialization of Adjacency Matrix
where $\mathbb{T}$ represents the set of all feasible edges connecting dataset label nodes to unified label nodes, $x_{\boldsymbol{t}}\in\{0,1\}$ indicates whether edge $\boldsymbol{t}\in\mathbb{T}$ is selected, with $x_{\boldsymbol{t}}=1$ indicating the presence of edge $\boldsymbol{t}$ and 0 indicating its absence. $\mathcal{S}_{k}$ denotes the dataset to which edge $\boldsymbol{t}$ belongs, $L_{k}$ represents the label space of $\mathcal{S}_{k}$, and $\boldsymbol{t}(c)$ is the label connected by edge $\boldsymbol{t}$. $\lambda|L|$ is the penalty term for the number of nodes, encouraging the optimizer to merge similar nodes to obtain a more compact unified label space. $\mathcal{L}_{c}\left(S_{c}^{k},\tilde{S}_{c}^{k}\right)$ is primarily used to evaluate the quality of label merging, where $S_{c}^{k}$ represents the output results on class $c$ using the original dataset segmentation head, and $\tilde{S}_{c}^{k}$ represents the output results on class $c$ using segmentation heads from other datasets participating in the merging.

### v3

@@ 4.3 Ablation Study
To further explore the ability of our GNNs to construct a unified label space, we compared it with four alternative methods. The first method concatenates all dataset label spaces into a single unified space $\bigcup_{i=1}^{D}L_{i}$. The second approach constructs a unified label space by clustering text features based on cosine similarity using the DBSCAN [7] algorithm. In the third method, we train a segmentation network using an initial adjacency matrix, as outlined in Appendix B, without incorporating GNN training. The final method involves an ablation experiment, removing the label description module to observe its impact. Experimental results, presented in Table 5, demonstrate that the label space constructed based on GNNs can better assist in the learning of segmentation models. Unlike the first approach, our method optimizes model capacity by focusing on label relationships rather than dataset recognition. Compared to the second approach, our method can differentiate between classes with identical names but differing levels of granularity. By leveraging label descriptions to enrich semantic context, our approach constructs a more refined and functional label space.
@@ 5 Conclusion
Limitations. Although our approach does not require manual relabeling efforts, it still relies on fully annotated datasets for training, in contrast to weakly-supervised and unsupervised methods. We aim to explore ways to integrate these alternative methods in future research. Errors in the fully automated construction of a unified label space do present some safety risks for autonomous driving tasks. Therefore, we recommend introducing a manual review mechanism to address these issues. Additionally, using ChatGPT may generate inaccurate label descriptions, which could affect the prediction of label relationships. Therefore, we aim to improve the accuracy of label descriptions by incorporating label descriptions provided by official datasets as prompts.
6 Acknowledgments
This work was supported in part by NSFC Project (62176061), Shanghai Municipal Science and Technology Major Project (2021SHZDZX0103). The computations in this research were performed using the CFFF platform of Fudan University.
@@ Appendix A Training Strategy
The training process is divided into three distinct stages: Multi-SegHead Training Stage, Alternating GNNs Training Stage and Alternating SegNet Training Stage, as shown in Algorithm 1. Initially, we commence with the Multi-SegHead Training Stage for 100k iterations. After completing Multi-SegHead Training Stage, we discard the multi-heads and initialize the UniSegHead for predicting class probabilities during subsequent training phases. Subsequently, we alternate between the Alternating GNNs Training Stage and Alternating SegNet Training Stage for a total of three cycles, each lasting 20k iterations. Notably, we prioritize the Alternating GNNs Training stage at the beginning of each cycle.
@@
$\mathcal{L}_{m}(\mathbf{y},\mathbf{s})=-\sum^{|L_{i}|}_{c=1}y_{c}\log(\operatorname{softmax}(\mathbf{W}_{i}\textbf{p})_{c}).$ (9)
@@ Appendix B Selection of Unified Label Node Quantity and Initialization of Adjacency Matrix
where $\mathbb{T}$ represents the set of all feasible edges connecting dataset label nodes to unified label nodes, $x_{\boldsymbol{t}}\in\{0,1\}$ indicates whether edge $\boldsymbol{t}\in\mathbb{T}$ is selected, with $x_{\boldsymbol{t}}=1$ indicating the presence of edge $\boldsymbol{t}$ and 0 indicating its absence. $\mathcal{S}_{k}$ denotes the dataset to which edge $\boldsymbol{t}$ belongs, $L_{k}$ represents the label space of $\mathcal{S}_{k}$, and $\boldsymbol{t}(c)$ is the label connected by edge $\boldsymbol{t}$. $\lambda|L|$ is the penalty term for the number of nodes, encouraging the optimizer to merge similar nodes to obtain a more compact unified label space. Based on experimental insights and referencing paper [52], we selected the hyperparameter $\lambda=0.5$. $\mathcal{L}_{c}\left(S_{c}^{k},\tilde{S}_{c}^{k}\right)$ is primarily used to evaluate the quality of label merging, where $S_{c}^{k}$ represents the output results on class $c$ using the original dataset segmentation head, and $\tilde{S}_{c}^{k}$ represents the output results on class $c$ using segmentation heads from other datasets participating in the merging.
@@ Appendix F Visualization
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The final paragraph of our introduction (section 1) provides a comprehensive summary of the paper’s contributions, which align with the claims made in the abstract.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: In the third paragraph of our conclusion (section 5), we discuss the limitations of our work, including factors influencing performance, and computational efficiency considerations.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: Our paper does not include theoretical results.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: In section 3, Appendix A, we provide detailed implementation steps, ensuring that our experiments can be replicated. Additionally, in section 4, we thoroughly describe the settings of our experiments, further facilitating reproducibility.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We provide reproducible code along with detailed instructions, ensuring that the main experimental results can be faithfully reproduced.
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: In section 4, we provide comprehensive details of our experimental setting.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [No]
  28. Justification: Training on numerous datasets is time-consuming, with each experiment taking over a week. During our experiments, we observed consistent and stable performance in terms of mIoU for our model.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: In section 4, we specify that our experiments were conducted on four 80G A100 GPUs.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We have read and adhered to the NeurIPS Code of Ethics in conducting our research.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [Yes]
  40. Justification: In the first paragraph of our conclusion (section 5), we thoroughly discuss the broader impacts of our work.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: Our research utilizes openly available datasets for training, and our models are primarily used for semantic segmentation tasks. Therefore, there are no such risks associated with our work.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: In section 4, we clearly specify the assets used in our work, including datasets and models. We provide proper citations to the original papers and ensure that the licenses and terms of use are explicitly mentioned and respected.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [Yes]
  52. Justification: We provide our code as an asset and include comprehensive documentation alongside it, covering details such as training procedures, licensing, limitations, and any necessary consent obtained from individuals whose assets are used.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: Our paper does not involve crowdsourcing experiments or research with human subjects.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: Our work does not involve human subjects.